A flywheel with a radius of 0.500m starts from rest and accelerates with a constant angular acceleration of 0.800rad/s2 .

a -Compute the magnitude of the resultant acceleration of a point on its rim after it has turned through 60.0 ?.

b-Compute the magnitude of the tangential acceleration of a point on its rim after it has turned through 120.0

c-Compute the magnitude of the radial acceleration of a point on its rim after it has turned through 120.0

?.?.
Answer #1

after 60o = ?/3 = 1.05 rad

angular velocity = sqrt(2*0.8*1.05) = 1.3 rad/sec

radial acc = (1.3*0.5)2/0.5 = 0.845 m/sec2

resultant acc = sqrt((0.8*0.5)2+ 0.8452)

= 0.934 m/sec2

tangential acc = 0.8*0.5 = 0.4m/sec2

after 120o = 2?/3 = 2.1 rad

angular velocity = sqrt(2*2.1*0.8) = 1.83 rad/sec

radial acc = (1.83*0.5)2/0.5 = 1.67m/sec2